HUBBLE SPACE TELESCOPE DAILY REPORT #937 PERIOD COVERED: 0800 EST Daily Status Report as of 208/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: The following proposals were not executed due to the current Software Sunpoint Safemode event: (See Operations Note 2.4) 1.1 Completed WFC 4599 (Optical Morphology of Radio Quiet, Cluster Galaxies at Redshift Z = 3.4) 1.2 Completed FOS/BL 5050 (Cycle 3 Internal Wavelenght Calibration) 1.3 Completed Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 1.4 Uplinked Updated Ephemeris Tables () 1.5 Completed FOS/RD 4772 (Black Holes, Stellar Dynamics and Populations in the Nuclei) 1.6 Completed WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 0 Successful : 0 Scheduled Reacquisitions : 0 Successful : 0 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 0 Successful : 0 2.3 Significant HSTARs: None 2.4 Operations Note: The HST entered software sunpoint safemode at approximately 207/0724Z ( Monday 26, 1993). The Spacecraft on-board computer detected a difference between the expected and actual values for the solar array wings during a scheduled wing maneuver of 30 degrees. Transition was successfully made to the Health and Safety load at 208/0600Z. The Safe Mode event occurred during a ZOE, and the engineering data was successfully played back through the MILA ground station. A test was successfully conducted to turn on the DCE and SADE units yesterday, and things looked okay. All data conditioned by these units were nominal. Functional testing is in planning. Investigations continues, with a possible test run on the spacecraft today or tomorrow to check overall SADE/SA operations. The return to the normal science SMS is still on hold, with a possible initiation of normal science operations this weekend. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: The following proposals were scheduled to be executed, but will be canceled due to the current Software Sunpoint Safemode event: (See Operations Note 2.4) 3.1 Complete FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) 3.2 Complete GHRS/2 5040 (GHRS Detector Calibrations - Cycle 3) 3.3 Complete Two Sets of WFC 4381 (HST Medium Deep Survey - Cycle 3) 3.4 Complete HSP/UV1 1091 (Ultraviolet Pulsations from X-Ray Pulsars) 3.5 Complete FOS/RD 5050 (Cycle 3 Internal Wavelenght Calibration) 3.6 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.7 Complete Two Sets of PC 4929 (WF/PC Cycle 2 :Improve Dark Calibration) 3.8 Complete FOS/RD 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) 3.9 Complete FOS/BL 4548 (Search foe Evidence of Active Ice Volcanism of Europa Cycle 3 - High) 3.10 Complete FOS/RD/BL 4340 (The Origin of the Blue Featureless Continuum in Seyfert 2 Nuclei) 3.11 Complete PC 3912 (Cores of Early - Type Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-